# Variational inequalities and smooth-fit principle for singular stochastic control problems in Hilbert spaces

### v1

@@ 3.2. Dynamic programming equation and viscosity solutions
Let $x\in H$, $\varphi\in\mathcal{X}$, $I\in\hat{\mathcal{I}_{0}}$, and $\tau\in\mathcal{T}$. Assume that
(3.16) $\mathsf{E}\left[\int_{0^{-}}^{\tau}e^{-\rho t}|X^{x,I}_{t}-x|_{H}|\langle X^{x,I}_{t}-x,\vartheta_{t}\rangle_{H}|\mathrm{d}\nu_{t}\right]<\infty.$
Then, the following inequality form of Dynkin’s formula holds true:
$\mathsf{E}\left[e^{-\rho\tau}|X^{x,I}_{\tau}-x|^{3}_{H}\right]$
$\leq$ $\ \mathsf{E}\left[\int_{0}^{\tau}e^{-\rho t}\left(3|\sigma\sigma^{*}|_{\mathcal{L}_{1}(H)}|X^{x,I}_{t}-x|_{H}-\rho|X^{x,I}_{t}-x|_{H}^{3}\right)\mathrm{d}t+\int_{0^{-}}^{\tau}e^{-\rho t}3|X^{x,I}_{t}-x|_{H}\langle X^{x,I}_{t}-x,\vartheta_{t}\rangle_{H}\mathrm{d}\nu_{t}\right].$
@@
$Dg(y)=3|y-x|_{H}(y-x),\ \ \ D^{2}g(y)=3|y-x|_{H}^{-1}(y-x)\otimes(y-x)+3|y-x|_{H}\cdot\textbf{Id}_{H}.$
@@
$\mbox{Tr}\left[\sigma\sigma^{*}D^{2}g(y)\right]\leq|D^{2}g(y)|_{\mathcal{L}(H)}\mbox{Tr}\left[\sigma\sigma^{*}\right]\leq 6|y-x|_{H}|\sigma\sigma^{*}|_{\mathcal{L}_{1}(H)}.$
[…]
The next result is the dynamic programming principle for the singular stochastic control problem (2.20). We are not going to provide a proof here, as this would very much follow the arguments used in the finite-dimensional settings (see [22], [38], and [49], among others). As a matter of fact, the key steps in the proof are based on the flow property of the solution to the controlled dynamics, which is, in the present setting, guaranteed by the semigroup property.
Proposition 3.5.
[Dynamic Programming Principle for $V$] Recall (3.1). For each $\tau\in\mathcal{T}$, we have
(3.17) $V(x)=\inf_{I:=(\vartheta,\nu)\in\hat{\mathcal{I}_{0}}}\mathsf{E}\left[\int_{0^{-}}^{\tau}e^{-\rho t}\left(G(X^{x,I}_{t})\mathrm{d}t+\mathrm{d}\nu_{t}\right)+e^{-\rho\tau}V(X_{\tau}^{x,I})\right],\quad x\in H.$
We are then ready to provide the definition of viscosity solution to (3.2) and to prove the viscosity property of $V$ as in (2.20).
@@
    We say that $v\in C(H)$ is a viscosity supersolution to (3.2) at $x\in H$ if, for every $\varphi\in\mathcal{X}$ such that $0=v(x)-\varphi(x)=\min(v-\varphi)$, one has
    $\max\Big{\{}(\rho-\mathcal{G})\varphi(x)-G(x),\ \sup_{\theta\in\Theta}\big{\{}-\langle D{\varphi}(x),\theta\rangle_{H}-1\big{\}}\Big{\}}\geq 0.$
@@
    We say that $v\in C(H)$ is a viscosity subsolution to (3.2) at $x\in H$ if, for every $\varphi\in\mathcal{X}$ such that $0=v(x)-\varphi(x)=\max(v-\varphi)$, one has
    $\max\Big{\{}(\rho-\mathcal{G})\varphi(x)-G(x),\ \sup_{\theta\in\Theta}\big{\{}-\langle D{\varphi}(x),\theta\rangle_{H}-1\big{\}}\Big{\}}\leq 0.$
@@
(Subsolution property.) Let $x\in H$, $\varphi\in\mathcal{X}$ be such that $0=V(x)-\varphi(x)=\max(V-\varphi)$.
@@
We now aim at taking limits as $h\to 0^{+}$ in (3.18). To that end, note that the limits in the right-hand side of (3.18) can be interchanged with the expectation: By the monotone convergence theorem for the integral term; because of the dominated convergence theorem for the third addend, since $V$ has sub-polynomial growth (cf. Proposition 3.1), and since $X^{x,I}_{h}=X^{x,0}_{h}+\zeta\theta$ (cf. (2.7)) and $\mathsf{E}[\sup_{t\in[0,T]}|X^{x,0}_{t}|^{p}]<\infty$, for $p\geq 2$ and for any $T>0$ by (2.11). Hence, due to $X^{x,I}_{h}\to x+\zeta\theta$ as $h\to 0^{+}$, we find
@@
$\varphi(x)\leq\zeta+\varphi(x+\zeta\theta),$
@@
$-\langle D\varphi(x),\theta\rangle_{H}\leq 1.$
@@
$\sup_{\theta\in\Theta}\left\{-\langle D\varphi(x),\theta\rangle_{H}-1\right\}\leq 0.$
@@
$\varphi(x)\leq\mathsf{E}\left[\int_{0}^{\tau_{R}\wedge h}e^{-\rho t}G(X^{x,0}_{t})\mathrm{d}t+e^{-\rho(\tau_{R}\wedge h)}\varphi(X_{\tau_{R}\wedge h}^{x,0})\right].$
Using now Proposition 3.3, we find
$\mathsf{E}\bigg{[}\int_{0}^{\tau_{R}\wedge h}e^{-\rho t}\Big{(}\big{[}(\rho-\mathcal{G})\varphi\big{]}(X_{t}^{x,0})-G(X^{x,0}_{t})\Big{)}\mathrm{d}t\bigg{]}\leq 0.$
@@
$(\rho-\mathcal{G})\varphi(x)-G(x)\leq 0.$
@@
(Supersolution property.) Let now $x\in H$, $\varphi\in\mathcal{X}$ be such that $0=V(x)-\varphi(x)=\min(V-\varphi)$. Assume, by contradiction that there exists $\eta>0$ such that
(3.20) $\sup_{\theta\in\Theta}\left\{-\langle D\varphi(x),\theta\rangle_{H}-1\right\}\leq-2\eta$
@@
(3.21) $(\rho-\mathcal{G})\varphi(x)-G(x)\leq-2\eta.$
By continuity, for a suitable $\varepsilon>0$,
(3.22) $\sup_{\theta\in\Theta}\left\{-\langle D\varphi(y),\theta\rangle_{H}-1\right\}\leq-\eta,\ \ \ \ \forall y\in B_{|\cdot|_{H}}(x,\varepsilon).$
@@
(3.23) $(\rho-\mathcal{G})\varphi(y)-G(y)\leq-\eta,\ \ \ \ \forall y\in B_{|\cdot|_{H}}(x,\varepsilon),$
where
$B_{|\cdot|_{H}}(x,\varepsilon):=\{y\in H:|y-x|_{H}\leq\varepsilon\}.$
Let now $I=(\vartheta,\nu)\in\hat{\mathcal{I}}_{0}$ be arbitrary but fixed, and set
@@
with the convention $\inf\emptyset=\infty$, which still provides a sense to the formulae below. In the following, we are going simply to write $\tau_{\varepsilon}$ instead of $\tau_{\varepsilon}^{I}$, unless it becomes important to stress the explicit dependence on $I$. By Assumption 2.4 and the fact that $\varphi\in\mathcal{X}$, we have for a constant $C>0$
$\sup_{B_{|\cdot|_{H}}(x,\varepsilon)}|D\varphi|_{L^{\infty}}\leq C\cdot\sup_{B_{|\cdot|_{H}}(x,\varepsilon)}|D\varphi|_{\mathcal{D}(\mathcal{A})}=:\widetilde{C}<\infty.$
Then, since $q\geq q_{o}\mathbf{1}$, we have
$\mathsf{E}\bigg{[}\int_{0^{-}}^{\tau_{\varepsilon}}e^{-\rho t}\left|\langle D\varphi(X^{x,I}_{t}),\vartheta_{t}\rangle_{H}\right|\mathrm{d}\nu_{t}\bigg{]}\leq\widetilde{C}\mathsf{E}\bigg{[}\int_{0^{-}}^{\tau_{\varepsilon}}e^{-\rho t}\langle\mathbf{1},\vartheta_{t}\rangle_{H}\mathrm{d}\nu_{t}\bigg{]}$
(3.24) $=\frac{\widetilde{C}}{q_{o}}\mathsf{E}\bigg{[}\int_{0^{-}}^{\infty}e^{-\rho t}\langle q_{o}\mathbf{1},\vartheta_{t}\rangle_{H}\mathrm{d}\nu_{t}\bigg{]}\leq\frac{\widetilde{C}}{q_{o}}\mathsf{E}\bigg{[}\int_{0^{-}}^{\infty}e^{-\rho t}\langle q,\vartheta_{t}\rangle_{H}\mathrm{d}\nu_{t}\bigg{]}$
$=\frac{\widetilde{C}}{q_{o}}\mathsf{E}\bigg{[}\int_{0^{-}}^{\infty}e^{-\rho t}\mathrm{d}\nu_{t}\bigg{]}<\infty.$
Similarly, one gets
(3.25) $\mathsf{E}\left[\int_{0^{-}}^{\tau_{\varepsilon}}e^{-\rho t}\left|\langle X^{x,I}_{t}-x,\vartheta_{t}\rangle_{H}\right|\mathrm{d}\nu_{t}\right]<\infty.$
Hence, employing (3.22), (3.23), and Propositions 3.3 and 3.4, whose requirements are met due to (3.2) and (3.25), we find
$V(x)-\mathsf{E}[e^{-\rho\tau_{\varepsilon}}V(X_{\tau_{\varepsilon}}^{x,I})]\leq{\varphi}(x)-\mathsf{E}[e^{-\rho\tau_{\varepsilon}}{\varphi}(X_{\tau_{\varepsilon}}^{x,I})]$
$=-\mathsf{E}\left[e^{-\rho\tau_{\varepsilon}}\big{|}X_{\tau_{\varepsilon}}^{x,I}-x\big{|}^{3}_{H}\right]+\mathsf{E}\left[e^{-\rho\tau_{\varepsilon}}\big{|}X_{\tau_{\varepsilon}}^{x,I}-x\big{|}^{3}_{H}\right]$
$+\mathsf{E}\left[\int_{0}^{\tau_{\varepsilon}}e^{-\rho t}(\rho-\mathcal{G}){\varphi}(X_{t}^{x,I})\mathrm{d}t-\int_{0^{-}}^{\tau_{\varepsilon}}e^{-\rho t}\langle D\varphi(X^{x,I}_{t}),\vartheta_{t}\rangle_{H}\mathrm{d}\nu_{t}\right]$
$\leq-\varepsilon^{3}\mathsf{E}\left[e^{-\rho\tau_{\varepsilon}}\mathds{1}_{\{\tau_{\varepsilon}<\infty\}}\right]+\mathsf{E}\left[\int_{0}^{\tau_{\varepsilon}}e^{-\rho t}3|\sigma\sigma^{*}|_{\mathcal{L}_{1}(H)}\varepsilon\mathrm{d}t+\int_{0^{-}}^{\tau_{\varepsilon}}3e^{-\rho t}\varepsilon^{2}\mathrm{d}\nu_{t}\right]$
$+\mathsf{E}\left[\int_{0^{-}}^{\tau_{\varepsilon}}e^{-\rho t}(G(X_{t}^{x,I})\mathrm{d}t+\mathrm{d}\nu_{t})\right]-\eta\mathsf{E}\left[\int_{0^{-}}^{\tau_{\varepsilon}}e^{-\rho t}(\mathrm{d}t+\mathrm{d}\nu_{t})\right].$
The latter implies
@@
$\leq-\varepsilon^{3}\mathsf{E}\left[e^{-\rho\tau_{\varepsilon}}\right]+\mathsf{E}\left[\int_{0}^{\tau_{\varepsilon}}e^{-\rho t}3|\sigma\sigma^{*}|_{\mathcal{L}_{1}(H)}\varepsilon\mathrm{d}t+\int_{0^{-}}^{\tau_{\varepsilon}}3e^{-\rho t}\varepsilon^{2}\mathrm{d}\nu_{t}\right]-\eta\mathsf{E}\left[\int_{0^{-}}^{\tau_{\varepsilon}}e^{-\rho t}(\mathrm{d}t+\mathrm{d}\nu_{t})\right],$
where we have used that $e^{-\rho\tau_{\varepsilon}}\mathds{1}_{\{\tau_{\varepsilon}=+\infty\}}=0$ by (3.6).
Stressing now the dependency of $\tau_{\varepsilon}$ with respect to the given and fixed $I\in\hat{\mathcal{I}}_{0}$, taking the supremum on both terms of (3.26) with respect to $I\in\hat{\mathcal{I}}_{0}$, using (3.17), and considering the identity
@@
we obtain
$0\leq\sup_{I\in\hat{\mathcal{I}}_{0}}\left\{-\varepsilon^{3}\mathsf{E}\left[\rho-\int_{0}^{\tau_{\varepsilon}^{I}}e^{-\rho t}\mathrm{d}t\right]+\mathsf{E}\left[\int_{0}^{\tau_{\varepsilon}^{I}}e^{-\rho t}\left(3|\sigma\sigma^{*}|_{\mathcal{L}_{1}(H)}\varepsilon-\eta\right)\mathrm{d}t+\int_{0^{-}}^{\tau_{\varepsilon}^{I}}e^{-\rho t}(3\varepsilon^{2}-\eta)\mathrm{d}\nu_{t}\right]\right\};$
@@
$0\leq-\rho\varepsilon^{3}+\sup_{I\in\hat{\mathcal{I}}_{0}}\mathsf{E}\left[\int_{0}^{\tau_{\varepsilon}^{I}}e^{-\rho t}\left(3\varepsilon|\sigma\sigma^{*}|_{\mathcal{L}_{1}(H)}+\varepsilon^{3}-\eta\right)\mathrm{d}t+\int_{0^{-}}^{\tau_{\varepsilon}^{I}}e^{-\rho t}(3\varepsilon^{2}-\eta)\mathrm{d}\nu_{t}\right],$
Taking $\varepsilon$ small enough with respect to $\eta$, which can be done without loss of generality, the integrands on the right-hand side of the last display formula are negative, so that the supremum is nonpositive. Hence, we reach a contradiction and the supersolution property of $V$ is proved. ∎
@@
It is worth emphasizing that the argument of the proof of the supersolution property can be clearly applied also in the finite-dimensional setting. As a consequence, this novel argument is able to reduce substantially the technicalities that are typically needed in $\mathbb{R}^{n}$, $n\geq 1$, in order to show the supersolution (respectively, subsolution) property in minimization problems (respectively, maximization problems) involving singular controls (see, e.g., [16], [34, Ch. VIII], [38], and [49], among others).

### v2

@@ 3.2.1. Dynkin’s formulae
Let $x\in H$, $z\in\mathcal{D}(\mathcal{A})$, $I\in\hat{\mathcal{I}}_{0}$. Let $\tau\in\mathcal{T}$ be bounded and such that there exists $M>0$ for which $\sup_{0\leq t\leq\tau}|X^{x,I}_{t}|\leq M$. Let $g_{0}\in C^{2}(\mathbb{R}^{+};\mathbb{R})$ be nondecreasing and such that $g^{\prime}_{0}(0)=0$, and let $g(y):=g_{0}(|y-z|_{H})$. Then, the following inequality form of Dynkin’s formula holds true:
$\mathsf{E}[e^{-\rho\tau}g(X^{x,I}_{\tau})]$ $\leq g(x)+\mathsf{E}\bigg[\int_{0}^{\tau}e^{-\rho t}\Big[\langle\mathcal{A}z,Dg(X_{t}^{x,I})\rangle_{H}+\frac{1}{2}\mathrm{Tr}[\sigma\sigma^{*}D^{2}g(X_{t}^{x,I})]-\rho g(X^{x,I}_{t})\Big]\mathrm{d}t\bigg]$
$+\mathsf{E}\bigg[\int_{0^{-}}^{\tau}e^{-\rho t}\langle Dg(X^{x,I}_{t}),\vartheta_{t}\rangle_{H}\mathrm{d}\nu_{t}\bigg].$
Proof.
As in the proof of Proposition 3.2, we get
$\mathsf{E}[e^{-\rho\tau}g(X^{n;x,I}_{\tau})]$
$=g(x)+\mathsf{E}\bigg[\int_{0}^{\tau}e^{-\rho t}\Big((\mathcal{G}_{n}-\rho)g\Big)(X^{n;x,I}_{t})\mathrm{d}t\bigg]+\mathsf{E}\bigg[\int_{0^{-}}^{\tau}e^{-\rho t}\langle Dg(X^{n;x,I}_{t}),\vartheta_{t}\rangle_{H}\mathrm{d}\nu_{t}\bigg]$
$=g(x)+\mathsf{E}\bigg[\int_{0}^{\tau}e^{-\rho t}\Big(\langle\mathcal{A}_{n}X_{t}^{n,x,I},Dg(X_{t}^{n,x,I})\rangle_{H}+\frac{1}{2}\mbox{Tr}[\sigma\sigma^{*}D^{2}g(X_{t}^{n,x,I})]-\rho g(X^{n;x,I}_{t})\Big)\mathrm{d}t\bigg]$
$\ \ \ +\mathsf{E}\bigg[\int_{0^{-}}^{\tau}e^{-\rho t}\langle Dg(X^{n;x,I}_{t}),\vartheta_{t}\rangle_{H}\mathrm{d}\nu_{t}\bigg].$
Considering (3.17), the fact that $g_{0}^{\prime}\geq 0$, and the dissipativity of the Yosida approximants $\mathcal{A}_{n}$ inherited by the dissipativity of $\mathcal{A}$, we get
$\langle\mathcal{A}_{n}X_{t}^{n,x,I},Dg(X_{t}^{n,x,I})\rangle_{H}$ $=\langle\mathcal{A}_{n}(X_{t}^{n,x,I}-z),Dg(X_{t}^{n,x,I})\rangle_{H}+\langle\mathcal{A}_{n}z,Dg(X_{t}^{n,x,I})\rangle_{H}$
$\leq\langle\mathcal{A}_{n}z,Dg(X_{t}^{n,x,I})\rangle_{H}.$
Hence,
$\mathsf{E}[e^{-\rho\tau}g(X^{n;x,I}_{\tau})]$
$\leq g(x)+\mathsf{E}\bigg[\int_{0}^{\tau}e^{-\rho t}\Big(\langle\mathcal{A}_{n}z,Dg(X_{t}^{n,x,I})\rangle_{H}+\frac{1}{2}\mbox{Tr}[\sigma\sigma^{*}D^{2}g(X_{t}^{n,x,I})]-\rho g(X^{n;x,I}_{t})\Big)\mathrm{d}t\bigg]$
$\ \ \ +\mathsf{E}\bigg[\int_{0^{-}}^{\tau}e^{-\rho t}\langle Dg(X^{n;x,I}_{t}),\vartheta_{t}\rangle_{H}\mathrm{d}\nu_{t}\bigg].$
Now, by the reverse Fatou’s lemma (thanks to integrability properties as those in the proof of Proposition 3.2) we can take limit superior as $n\to\infty$ in order to obtain the claim. ∎
The following corollary will be exploited in the proof of the viscosity supersolution property of $V$ (see Theorem 3.8 below).
Corollary 3.4.
Let $x\in H$, $z\in\mathcal{D}(\mathcal{A})$, $I\in\hat{\mathcal{I}}_{0}$. Let $\tau\in\mathcal{T}$ be bounded and assume that, for some $\varepsilon>0$, it holds $|X^{x,I}_{t}-z|_{H}\leq\varepsilon$ for all $t\in[0,\tau].$ Then, there exists a constant $c_{o}>0$ such that
(3.18) $\mathsf{E}\left[e^{-\rho\tau}|X^{x,I}_{\tau}-z|^{3}_{H}\right]\leq c_{o}(\varepsilon+\varepsilon^{2})\,\mathsf{E}\left[\int_{0^{-}}^{\tau}e^{-\rho t}(\mathrm{d}t+\mathrm{d}\nu_{t})\right].$
@@
$Dg(y)=3|y-z|_{H}(y-z),\ \ \ D^{2}g(y)=3|y-z|_{H}^{-1}(y-z)\otimes(y-z)+3|y-z|_{H}\cdot\textbf{Id}_{H}.$
@@
$\mbox{Tr}\left[\sigma\sigma^{*}D^{2}g(y)\right]\leq|D^{2}g(y)|_{\mathcal{L}(H)}\mbox{Tr}\left[\sigma\sigma^{*}\right]\leq 6|y-z|_{H}|\sigma\sigma^{*}|_{\mathcal{L}_{1}(H)}.$
Then, the claim follows from Proposition 3.3. ∎
3.2.2. Dynamic Programming Principle
The next result is the dynamic programming principle for the singular stochastic control problem (2.22). We provide a sketch of its proof in Appendix A.
Proposition 3.5 (Dynamic Programming Principle for $V$).
Recall (3.1). We have
(3.19) $V(x)=\inf_{(I,\tau^{I})\in\hat{\mathcal{I}_{0}}\times\mathcal{T}}\mathsf{E}\left[\int_{0^{-}}^{\tau^{I}}e^{-\rho t}\left(G(X^{x,I}_{t})\mathrm{d}t+\mathrm{d}\nu_{t}\right)+e^{-\rho\tau^{I}}V(X_{\tau^{I}}^{x,I})\right],\quad x\in H.$
3.2.3. The value function as viscosity solution of the variational inequality
We are ready to provide the definition of viscosity solution to (3.3) and to prove that $V$ defined in (2.22) is actually a viscosity solution to the variational inequality.
@@ 3.2.3. The value function as viscosity solution of the variational inequality
    We say that $v\in C(H)$ is a viscosity supersolution to (3.3) at $x\in H$ if, for every $\psi=\varphi+g$ with $\varphi\in\mathcal{X}$, $g\in\mathcal{R}^{0}$ such that $0=v(x)-\psi(x)=\min(v-\psi)$, one has max{ρψ(x) -12Tr[σσ^* D^2ψ(x)]-x,ADφ(x)_H-G(x), supΘ{- Dψ(x),θ_H-1} }0.
@@
    We say that $v\in C(H)$ is a viscosity subsolution to (3.3) at $x\in H$ if, for every $\psi=\varphi-g$ with $\varphi\in\mathcal{X}$, $g\in\mathcal{R}^{0}$ such that $0=v(x)-\psi(x)=\max(v-\psi)$, one has max{ρψ(x) -12Tr[σσ^* D^2ψ(x)]-x,ADφ(x)_H-G(x), supΘ{- Dψ(x),θ_H-1} }0.
@@
Remark 3.7.
In the context of classical (regular) stochastic control, the definition of viscosity solution based on the set of test functions $\mathcal{X}$ combined with the set of radial functions $\mathcal{R}^{0}$ is typically sufficient to prove comparison results for viscosity solutions of the PDE (see [28, Ch. 3]); hence, to characterize the value function of the control problem as the unique solution to the associated dynamic programming equation.
In this paper, we do not address the relevant topic of uniqueness (see Remark 3.10 below), leaving it as a future research topic. Instead, we will use test functions from the set $\mathcal{X}$ of quadratic form in Section 4 in order to prove our regularity result (see, in particular, the proof of Proposition 4.17).
@@
(Subsolution property.) Let $x\in H$ and $\psi=\varphi+g$ with $\varphi\in\mathcal{X}$, $g\in\mathcal{R}^{0}$ such that $0=V(x)-\psi(x)=\max(V-\psi)$.
@@
We now aim at taking limits as $h\to 0^{+}$ in (3.20). To that end, note that the limits in the right-hand side of (3.20) can be interchanged with the expectation: By the monotone convergence theorem for the integral term; because of the dominated convergence theorem for the third addend, since $V$ has sub-quadratic growth (cf. Proposition 3.1), and since $X^{x,I}_{h}=X^{x,0}_{h}+\zeta\theta$ (cf. (2.7)) and $\mathsf{E}[\sup_{t\in[0,T]}|X^{x,0}_{t}|^{2}]<\infty$, for any $T>0$, by (2.12). Hence, due to $X^{x,I}_{h}\to x+\zeta\theta$ as $h\to 0^{+}$, we find
@@
$\psi(x)\leq\zeta+\psi(x+\zeta\theta),$
@@
$\langle D\psi(x),\theta\rangle_{H}\leq 1.$
@@
$\sup_{\theta\in\Theta}\left\{-\langle D\psi(x),\theta\rangle_{H}-1\right\}\leq 0.$
@@
$\psi(x)\leq\mathsf{E}\left[\int_{0}^{\tau_{R}\wedge h}e^{-\rho t}G(X^{x,0}_{t})\mathrm{d}t+e^{-\rho(\tau_{R}\wedge h)}\psi(X_{\tau_{R}\wedge h}^{x,0})\right].$
Using now Propositions 3.2 and 3.3, we find
$\mathsf{E}\bigg[\int_{0}^{\tau_{R}\wedge h}e^{-\rho t}\Big(\rho\psi(X_{t}^{x,0})-\frac{1}{2}\mathrm{Tr}[\sigma\sigma^{*}D^{2}\psi(X_{t}^{x,0})]-\langle X_{t}^{x,0},\mathcal{A}D\varphi(X_{t}^{x,0})\rangle-G(X^{x,0}_{t})\Big)\mathrm{d}t\bigg]\leq 0.$
@@
$\rho\psi(x)-\frac{1}{2}\mathrm{Tr}[\sigma\sigma^{*}D^{2}\psi(x)]-\langle x,\mathcal{A}D\varphi(x)\rangle_{H}-G(x)\leq 0$
@@
(Supersolution property.) Let now $x\in H$, $\psi=\varphi-g$, with $\varphi\in\mathcal{X}$ and $g\in\mathcal{R}^{0}$ be such that $0=V(x)-\psi(x)=\min(V-\psi)$. Assume, by contradiction that there exists $\eta>0$ such that
(3.22) $\sup_{\theta\in\Theta}\left\{-\langle D\psi(x),\theta\rangle_{H}-1\right\}\leq-2\eta$
@@
(3.23) $\rho\psi(x)-\frac{1}{2}\mathrm{Tr}[\sigma\sigma^{*}D^{2}\psi(x)]-\langle x,\mathcal{A}D\varphi(x)\rangle_{H}-G(x)\leq-2\eta.$
By continuity, for a suitable $\varepsilon_{o}>0$,
(3.24) $\sup_{\theta\in\Theta}\left\{-\langle D\psi(y),\theta\rangle_{H}-1\right\}\leq-\eta,\ \ \ \ \forall y\in B_{|\cdot|_{H}}(x,\varepsilon_{o}).$
@@
(3.25) $\rho\psi(y)-\frac{1}{2}\mathrm{Tr}[\sigma\sigma^{*}D^{2}\psi(y)]-\langle y,\mathcal{A}D\varphi(y)\rangle_{H}-G(y)\leq-\eta,\ \ \ \ \forall y\in B_{|\cdot|_{H}}(x,\varepsilon_{o}),$
where $B_{|\cdot|_{H}}(x,\varepsilon_{o}):=\{y\in H:|y-x|_{H}\leq\varepsilon_{o}\}.$
Let now $I=(\vartheta,\nu)\in\hat{\mathcal{I}}_{0}$ be arbitrary but fixed, and set, for $\varepsilon\in(0,\varepsilon_{o})$,
@@
with the convention $\inf\emptyset=\infty$, which still provides a sense to the formulae below. In the following, we are going simply to write $\tau_{\varepsilon}$ instead of $\tau_{\varepsilon}^{I}$, unless it becomes important to stress the explicit dependence on $I$.
Now pick $z\in\mathcal{D}(\mathcal{A})$ such that $|z-x|<\varepsilon/2$, which is possible as $\mathcal{D}(\mathcal{A})$ is dense in $H$. Employing (3.24), (3.25), and Propositions 3.2-3.3 and Corollary 3.4, we find for each $T>0$
$V(x)-\mathsf{E}[e^{-\rho{(\tau_{\varepsilon}\wedge T)}}V(X_{{\tau_{\varepsilon}\wedge T}}^{x,I})]\leq{\psi}(x)-\mathsf{E}[e^{-\rho{(\tau_{\varepsilon}\wedge T)}}{\psi}(X_{\tau_{\varepsilon}}^{x,I})]$
$=-\mathsf{E}\left[e^{-\rho{(\tau_{\varepsilon}\wedge T)}}\big|X_{{\tau_{\varepsilon}\wedge T}}^{x,I}-z\big|^{3}_{H}\right]+\mathsf{E}\left[e^{-\rho{(\tau_{\varepsilon}\wedge T)}}\big|X_{{\tau_{\varepsilon}\wedge T}}^{x,I}-z\big|^{3}_{H}\right]+{\psi}(x)-\mathsf{E}[e^{-\rho{(\tau_{\varepsilon}\wedge T)}}{\psi}(X_{\tau_{\varepsilon}}^{x,I})]$
$\leq-\frac{\varepsilon^{3}}{8}\mathsf{E}\left[e^{-\rho{(\tau_{\varepsilon}\wedge T})}\mathds{1}_{{{\{\tau_{\varepsilon}<T\}}}}\right]+\mathsf{E}\left[c_{o}(\varepsilon+\varepsilon^{2})\int_{0^{-}}^{\tau}e^{-\rho t}(\mathrm{d}t+\mathrm{d}\nu_{t})\right]$
$+\mathsf{E}\left[\int_{0}^{{\tau_{\varepsilon}\wedge T}}e^{-\rho t}\Big(\rho\psi(X_{t}^{x,I})-\frac{1}{2}\mathrm{Tr}[\sigma\sigma^{*}D^{2}\psi(X^{x,I}_{t})]-\langle X^{x,I}_{t},\mathcal{A}D\varphi(X^{x,I}_{t})\rangle_{H}\Big)\mathrm{d}t\right]$
$-\mathsf{E}\left[\int_{0^{-}}^{\tau_{\varepsilon}\wedge T}e^{-\rho t}\langle D\psi(X^{x,I}_{t}),\vartheta_{t}\rangle_{H}\mathrm{d}\nu_{t}\right]$
$\leq-\frac{\varepsilon^{3}}{8}\mathsf{E}\left[e^{-\rho{(\tau_{\varepsilon}\wedge T})}\mathds{1}_{{{\{\tau_{\varepsilon}<T\}}}}\right]+\mathsf{E}\left[c_{o}(\varepsilon+\varepsilon^{2})\int_{0^{-}}^{\tau}e^{-\rho t}(\mathrm{d}t+\mathrm{d}\nu_{t})\right]$
$+\mathsf{E}\left[\int_{0^{-}}^{{\tau_{\varepsilon}\wedge T}}e^{-\rho t}(G(X_{t}^{x,I})\mathrm{d}t+\mathrm{d}\nu_{t})\right]-\eta\mathsf{E}\left[\int_{0^{-}}^{{\tau_{\varepsilon}\wedge T}}e^{-\rho t}(\mathrm{d}t+\mathrm{d}\nu_{t})\right].$
Letting $T\uparrow\infty$ in the above inequality and using the fact that $e^{-\rho\tau_{\varepsilon}}\mathds{1}_{\{\tau_{\varepsilon}=+\infty\}}=0$ by (3.7), we obtain
@@
$\leq-\frac{\varepsilon^{3}}{8}\mathsf{E}\left[e^{-\rho\tau_{\varepsilon}}\right]+\left(c_{o}(\varepsilon+\varepsilon^{2})-\eta\right)\mathsf{E}\left[\int_{0^{-}}^{\tau_{\varepsilon}}e^{-\rho t}(\mathrm{d}t+\mathrm{d}\nu_{t})\right],$
The above inequality was obtained for a fixed $I\in\hat{\mathcal{I}}_{0}$. Now, stressing again the dependency of $\tau_{\varepsilon}$ with respect to $I\in\hat{\mathcal{I}}_{0}$, we take the supremum on both terms of (3.26) with respect to $I\in\hat{\mathcal{I}}_{0}$, we use (3.19), and consider the identity
@@
to obtain
$0\leq\sup_{I\in\hat{\mathcal{I}}_{0}}\left\{-\frac{\varepsilon^{3}}{8}\mathsf{E}\left[{1-\rho}\int_{0}^{\tau_{\varepsilon}^{I}}e^{-\rho t}\mathrm{d}t\right]+\left(c_{o}(\varepsilon+\varepsilon^{2})-\eta\right)\mathsf{E}\left[\int_{0^{-}}^{\tau_{\varepsilon}^{I}}e^{-\rho t}(\mathrm{d}t+\mathrm{d}\nu_{t})\right]\right\};$
@@
$0\leq-\frac{\varepsilon^{3}}{8}+\sup_{I\in\hat{\mathcal{I}}_{0}}\mathsf{E}\left[\left(c_{o}(\varepsilon+\varepsilon^{2})+\frac{\rho\varepsilon^{3}}{8}-\eta\right)\int_{0^{-}}^{\tau_{\varepsilon}^{I}}e^{-\rho t}(\mathrm{d}t+\mathrm{d}\nu_{t})\right],$
Taking $\varepsilon$ small enough with respect to $\eta$, which can be done without loss of generality, the supremum is nonpositive. Hence, we reach a contradiction and the supersolution property of $V$ is proved. ∎
@@
It is worth emphasizing that the argument of the proof of the supersolution property can be clearly applied also in the finite-dimensional setting. As a consequence, this novel argument is able to reduce substantially the technicalities that are typically needed in $\mathbb{R}^{n}$, $n\geq 1$, in order to show the supersolution (respectively, subsolution) property in minimization problems (respectively, maximization problems) involving singular controls (see, e.g., [16], [35, Ch. VIII], [39], and [51], among others).